OpenStudy (anonymous):

f(x) = 4x^2 -3, find f(a+3) , show all work. plz !

OpenStudy (anonymous):

This is algebra?

Parth (parthkohli):

I see. Replace \(x \) with \(a + 3\) everywhere. \[ \color{Black}{\Rightarrow f(a + 3) = 4(a + 3)^2 - 3}\]I leave upon you simplifying that.

OpenStudy (anonymous):

pre- cal

OpenStudy (anonymous):

\[f(a+3)=4(a+3)^2-3=4(a^2+6a+9)-3=4a^2+24a+36-3=4a^2+24a+33\]

OpenStudy (anonymous):

so the final answer is what ? cz i cant see it

Parth (parthkohli):

Uhm, please don't ask for answers here.

OpenStudy (anonymous):

\[4a^2+24a+33\]

Parth (parthkohli):

Not allowed to answer/ask for them.

OpenStudy (anonymous):

lol

Parth (parthkohli):

What do you think is \(4(a + 3)^2 - 3\)?